# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 4.3 Small $\tau_{H}$ limit
$\omega_{h}\simeq-\frac{iT^{2}\sigma^{\omega}_{A}}{6\tau_{H}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}.$ (113)
The angular frequencies for the other two(axial) modes, $\omega^{\pm}_{{\mathfrak{a}}}$, are finite when the helical mode freezes, $\tau_{H}\rightarrow 0$, satisfying
$\omega^{\pm}_{{\mathfrak{a}}}\simeq\frac{T^{2}}{12\sigma^{\omega}_{A}\tau_{A}}\left(-i\pm\sqrt{\frac{36A^{2}\sigma^{\omega}_{A}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)}\tau_{A}^{2}\kappa_{\Omega}^{2}-1}\right)\,.$ (114)
The above expression generalizes Eq. (71), giving the energy dispersion relation for the CVW to the case when the axial charge is not conserved. As already noted in Sec. 3.2, the vector and axial fluctuations do not propagate at vanishing chemical potential. At finite chemical potential and finite (i.e., non-infinite) $\tau_{A}$, they propagate provided that $k$ exceeds some critical value. Using the large temperature expansion 32, we find that the critical value behaves as $\tau_{A}\kappa_{*}=\frac{7\pi^{2}}{12\alpha_{V}}$ for small chemical potential. In contrast, it finds a minimum in the limit of infinite chemical potential: $\tau_{A}\kappa_{*}^{\rm min}=k_{*}^{\rm min}\tau_{A}\Omega/T=\pi/4\sqrt{3}\simeq 0.45$, which coincides with the large chemical potential limit of Eq. (109) giving the propagation threshold in the case when $\tau_{A}\rightarrow 0$. Since the helical and axial degrees of freedom correspond to the same physical quantity at infinite vector chemical potential, it is reasonable that freezing of either the helical or the axial chemical potentials gives the same result in this degenerate limit.
@@ 4.4 Large wavenumber limit
$\omega^{\pm}_{h}=\pm kc_{h}(\alpha_{V})+O(k^{0})\,,\qquad c_{h}(\alpha_{V})=\frac{\Omega T}{2}\sqrt{\frac{\sigma^{\omega}_{A}(A^{2}+B^{2})-2AB\sigma^{\omega}_{H}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}}\,,$ (115)
where $c_{h}(\alpha_{V})$ is the velocity of the helical vortical wave in an unpolarized fluid with vector imbalance $\mu_{V}=\alpha_{V}T$, in the absence of helicity- or chirality-violating interactions, as obtained in Eq. (4.20) of Ref. HVWp1 . The axial mode turns out to be purely dissipative:
@@
At large wavenumbers, the purely imaginary phase velocity of the axial mode will settle to a constant negative value, which will depend on the dimensionless vector chemical potential $\alpha_{V}=\mu_{V}/T$ and on the axial and helicity relaxation times. These results can be checked in Figure 9.
@@ 4.5 Large chemical potential limit
In order to study the large chemical potential limit, we solve the eigenvalue equation Eq. (37) perturbatively in $T/|\mu_{V}|$. Using the expansion in Eq. (33), we can compute the energy dispersion relation as a Taylor series in $e^{-|\mu_{V}|/T}$. Expanding
$\omega=\omega_{-1}e^{|\mu_{V}|/T}+\omega_{0}+O(e^{-|\mu_{V}|/T})\,,$ (117)
@@
$\dfrac{12i\sigma_{A}\left(\sigma_{A}-\frac{1}{3}T^{2}\Delta H\right)\left(\pi^{2}(\tau_{A}^{-1}+\tau_{H}^{-1})-24i\omega_{-1}\right)\omega_{-1}^{2}}{9\pi^{2}T^{4}}e^{2|\mu_{V}|/T}+O(e^{|\mu_{V}|/T})=0\,.$ (118)
@@
$\omega_{-1}^{\pm}=0\hskip 14.22636pt\textrm{(double root)}\,,\hskip 28.45274pt\omega_{-1}^{\textrm{im.}}=-\dfrac{i}{24\pi^{2}}\left(\frac{1}{\tau_{H}}+\frac{1}{\tau_{A}}\right)\,.$ (119)
The contributions $\omega_{0}^{\pm}$ to modes with $\omega_{-1}^{\pm}$ come from the next-to-leading order contribution to Eq. (37), namely
$\tau_{A}^{-1}\tau_{H}^{-1}\left(9A^{2}\kappa_{\Omega}^{2}(\tau_{A}+\tau_{H})-2H\omega_{0}^{\pm}\left(\sigma_{A}-\frac{T^{2}}{3}\Delta H\right)\left(iT^{2}+6\sigma_{A}(\tau_{A}+\tau_{H})\omega_{0}^{\pm}\right)\right)=0.$ (120)
Expanding $\omega_{0}$ in powers of $T/|\mu_{V}|$ gives, to leading order in $T/|\mu_{V}|$:
$\omega_{\pm}\simeq\dfrac{-i\pi^{2}}{6(\tau_{A}+\tau_{H})}\left(1\pm\sqrt{1-\frac{48}{\pi^{2}}(\tau_{A}+\tau_{H})^{2}\kappa_{\Omega}^{2}}\right)\dfrac{T^{2}}{|\mu_{V}|^{2}}.$ (121)
Note that when both relaxation times are finite, there is no propagation for small wavenumbers. This is in contrast to the result obtained in Subsec. 3.3, where we considered $\tau_{A}=\infty$, and the analogous modes propagate for all wavelengths. Notably, the $\pm$ modes are sensitive to an emergent effective relaxation time $\tau_{eff}=\tau_{A}+\tau_{H}$ which can be traced to the interdependence of the axial and helical degrees of freedom in the degenerate limit. From Eq. (121), we extract the critical wavenumber $k_{*}$ above which these two modes start propagating:
$k_{*}=\dfrac{\kappa_{\Omega}^{*}T}{\Omega\tau_{H}}=\frac{\pi}{4\sqrt{3}}\frac{T}{(\tau_{A}+\tau_{H})\Omega}\,.$ (122)
Refer to caption Refer to caption
@@
$\delta\mu^{*}_{\chi}$ $=\dfrac{\frac{2\omega_{*}}{T^{2}}(A+s_{V}B)(\sigma^{\omega}_{A}-s_{V}\sigma^{\omega}_{H})+\frac{i}{3}\left(\frac{A}{\tau_{H}}+\frac{s_{V}B}{\tau_{A}}\right)}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}}\kappa_{\Omega}\delta\mu^{*}_{V},$
$\delta\mu^{*}_{\tilde{\chi}}$ $=\dfrac{\frac{2\omega_{*}}{T^{2}}(A-s_{V}B)(\sigma^{\omega}_{A}+s_{V}\sigma^{\omega}_{H})+\frac{i}{3}\left(\frac{A}{\tau_{H}}-\frac{s_{V}B}{\tau_{A}}\right)}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}}\kappa_{\Omega}\delta\mu^{*}_{V}.$ (123)
@@
$\delta\mu_{\chi}^{\pm}$ $\simeq\frac{AT^{2}}{4\sigma_{A}^{\omega}}\frac{\kappa_{\Omega}(\tau_{A}+\tau_{H})}{\omega_{\pm}(\tau_{A}+\tau_{H})+\frac{iT^{2}}{6\sigma^{\omega}_{A}}}\delta\mu_{V}^{\pm}\simeq-\frac{\kappa_{\Omega}(\tau_{A}+\tau_{H})}{3\alpha_{V}^{2}\omega_{\mp}(\tau_{A}+\tau_{H})}\dfrac{|\mu_{V}|}{T}\delta\mu_{V}^{\pm},$
$\delta\mu_{\tilde{\chi}}^{\pm}$ $\simeq\frac{AT^{2}}{4\sigma_{A}^{\omega}}\frac{\kappa_{\Omega}(\tau_{A}-\tau_{H})}{\omega_{\pm}(\tau_{A}+\tau_{H})+\frac{iT^{2}}{6\sigma_{A}}}\delta\mu_{V}^{\pm}\simeq-\dfrac{\kappa_{\Omega}(\tau_{A}-\tau_{H})}{3\alpha_{V}^{2}\omega_{\mp}(\tau_{A}+\tau_{H})}\dfrac{|\mu_{V}|}{T}\delta\mu_{V}^{\pm}.$ (124)
As in the case with $\tau_{A}\rightarrow\infty$ considered in Eq. (81), the amplitudes of $\delta\mu^{\pm}_{\chi}$ and $\delta\mu^{\pm}_{\tilde{\chi}}$ are $|\alpha_{V}|=|\mu_{V}|/T$ times larger than $\delta\mu_{V}^{\pm}$, due to the behaviour $\omega_{\mp}\sim\alpha_{V}^{-2}$. Contrary to Eq. (81), the degeneracy between $\delta\mu^{\pm}_{\chi}$ and $\delta\mu^{\pm}_{\tilde{\chi}}$ is lifted whenever both $\tau_{A}$ and $\tau_{H}$ are finite (i.e., non-zero and non-infinite).
@@
$\delta\mu_{\chi}^{\textrm{im.}}\simeq-\dfrac{2(\tau_{A}-\tau_{H})}{(\tau_{A}+\tau_{H})\alpha_{V}^{2}}e^{-|\alpha_{V}|}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}\,,\quad\delta\mu_{V}^{\textrm{im.}}\simeq-\dfrac{72i\kappa_{\Omega}\tau_{A}\tau_{H}}{\pi^{2}(\tau_{A}+\tau_{H})\alpha_{V}^{2}}e^{-2|\alpha_{V}|}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}$ (125)
We present in Fig. 7 the propagation properties of an initial Gaussian fluctuation in the vector chemical potential, given in Eq. (61), similar to the one studied in Fig. 3 (see the end of Sec. 3.3 for details). As opposed to Fig. 3, here we take a finite axial relaxation time, which we set to $\tau_{A}=\tau_{H}/2$. While the propagation of $\delta\bar{\mu}_{V}(t,z)$ and of $\delta\bar{\mu}_{\chi}(t,z)$ remains unchanged, the amplitude of $\delta\bar{\mu}_{\tilde{\chi}}(t,z)$ is reduced by a factor of about $3$ compared to the case when the axial charge is conserved.
@@ 4.6 Generic properties of spectra
Figure 8: Solution to Eq. (131) showing the values of the parameters (blue curve) for which the spectrum of excitations contains a merging point. The curve divides the parameter space into two regions: Above the curve, the spectrum has only one branch of propagating modes; Below the curve, the spectrum displays two disconnected branches of propagating modes.
@@
In Sec. 3.4, we have seen that there exists a value of the vector chemical potential which divides the spectrum of excitations into two regions, depending on whether there exist one or two branches of propagating modes (c.f. Fig. 5). The chemical potential separating both regions was found to be $\mu_{V}=1.4471\,T$, and the wavenumber at which the two propagating branches merge $\tau_{H}\kappa_{\Omega}=3.0642$. In this section, the presence of a finite axial relaxation time $\tau_{A}$ provides an extra parameter, and we expect to find a one-parameter family of such points. We will impose the existence of the merging point in order to find the curve separating both regions in terms of $\mu_{V}$ and $\tau_{H}/\tau_{A}\,$.
@@
$a(\alpha_{V})\omega^{3}+i\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)b(\alpha_{V})\omega^{2}-\left(c(\alpha_{V})\kappa_{\Omega}^{2}+\dfrac{\varsigma(\alpha_{V})}{\tau_{A}\tau_{H}}\right)\omega-i\kappa^{2}\delta(\alpha_{V},\tau_{A},\tau_{H})=0\,,$ (126)
where $a\,,b$ and $c$ are already given in (3.3), while $\varsigma$ and $\delta$ are defined as
$\varsigma=\dfrac{2}{9T^{2}}\left(\sigma_{A}^{\omega}-\frac{1}{3}T^{2}\Delta H\right)\,,\qquad\delta=\dfrac{A^{2}\tau_{A}+B^{2}\tau_{H}}{3H\tau_{A}\tau_{H}}\,.$ (127)
@@
$3a\omega_{\textrm{merg.}}^{2}+2ib\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)\omega_{\textrm{merg.}}-\left(c(\alpha_{V})\kappa_{\textrm{merg.}}^{2}+\dfrac{\varsigma(\alpha_{V})}{\tau_{A}\tau_{H}}\right)=0\,,$
@@
$\omega_{\textrm{merg.}}=-\dfrac{2ib}{6a}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)\,\qquad\kappa_{\textrm{merg.}}^{2}=\dfrac{b^{2}}{3ac}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)^{2}-\dfrac{\varsigma}{c}\frac{1}{\tau_{A}\tau_{H}}\,.$ (129)

### v2

@@ 4.3 Small $\tau_{H}$ limit
$\omega_{h}\simeq-\frac{iT^{2}\sigma^{\omega}_{A}}{6\tau_{H}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]},$ (109)
which agrees with Eq. (62). The angular frequencies for the other two(axial) modes, $\omega^{\pm}_{{\mathfrak{a}}}$, are finite when the helical mode freezes, $\tau_{H}\rightarrow 0$, satisfying
$\omega^{\pm}_{{\mathfrak{a}}}\simeq\frac{T^{2}}{12\sigma^{\omega}_{A}\tau_{A}}\\ \times\left(-i\pm\sqrt{\frac{36A^{2}\sigma^{\omega}_{A}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)}\tau_{A}^{2}\kappa_{\Omega}^{2}-1}\right)\,.$ (110)
The above expression generalizes Eq. (67), giving the energy dispersion relation for the CVW to the case when the axial charge is not conserved. As already noted in Sec. 3.2, the vector and axial fluctuations do not propagate at vanishing chemical potential. At finite chemical potential and finite (i.e., non-infinite) $\tau_{A}$, they propagate provided that $k$ exceeds some critical value. Using the large temperature expansion (29), we find that the critical value behaves as $\tau_{A}\kappa_{*}=7\pi^{2}/(12\alpha_{V})$ for small chemical potential. In contrast, it finds a minimum in the limit of infinite chemical potential: $\tau_{A}\kappa_{*}^{\rm min}=k_{*}^{\rm min}\tau_{A}\Omega/T=\pi/4\sqrt{3}\simeq 0.45$, which coincides with the large chemical potential limit of Eq. (106) giving the propagation threshold in the case when $\tau_{A}\rightarrow 0$. Since the helical and axial degrees of freedom correspond to the same physical quantity at infinite vector chemical potential, it is reasonable that freezing of either the helical or the axial chemical potentials gives the same result in this degenerate limit.
@@ 4.4 Large wavenumber limit
$\omega^{\pm}_{h}$ =±$=\pm kc_{h}(\alpha_{V})+O(k^{0})\,,$
$c_{h}(\alpha_{V})$ $=\frac{\Omega T}{2}\sqrt{\frac{\sigma^{\omega}_{A}(A^{2}+B^{2})-2AB\sigma^{\omega}_{H}}{H(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]}}\,,$ (111)
where $c_{h}(\alpha_{V})$ is the velocity of the helical vortical wave in an unpolarized fluid with vector imbalance $\mu_{V}=\alpha_{V}T$, in the absence of helicity- or chirality-violating interactions, as obtained in Eq. (123) of Ref. Morales-Tejera:2024uzg . The axial mode turns out to be purely dissipative:
@@
At large wavenumbers, the phase velocity of the axial mode is imaginary and its imaginary part will settle to a constant negative value, which depends on the dimensionless vector chemical potential $\alpha_{V}=\mu_{V}/T$ and on the axial and helicity relaxation times. These results will be verified in Figure 9.
@@ 4.5 Large chemical potential limit
In order to study the large chemical potential limit, we solve the eigenvalue equation Eq. (33) perturbatively in $T/\lvert\mu_{V}\rvert$. Using the expansion in Eq. (30), we can compute the energy dispersion relation as a Taylor series in $e^{-\lvert\mu_{V}\rvert/T}$. Expanding
$\omega=\omega_{-1}e^{\lvert\mu_{V}\rvert/T}+\omega_{0}+O(e^{-\lvert\mu_{V}\rvert/T})\,,$ (113)
@@
$\dfrac{4i\sigma_{A}}{3T^{4}}\left(\sigma_{A}-\frac{1}{3}T^{2}\Delta H\right)\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}-\frac{24i}{\pi^{2}}\omega_{-1}\right)\\ \times e^{2\lvert\mu_{V}\rvert/T}\omega_{-1}^{2}+O(e^{\lvert\mu_{V}\rvert/T})=0\,.$ (114)
@@
[…]
$\omega_{-1}^{\textrm{im.}}=-\dfrac{i\pi^{2}}{24}\left(\frac{1}{\tau_{H}}+\frac{1}{\tau_{A}}\right)\,.$ (115)
The contributions $\omega_{0}^{\pm}$ to modes with $\omega_{-1}^{\pm}=0$ come from the next-to-leading order contribution to Eq. (33), namely
$-\frac{iA^{2}\kappa_{\Omega}^{2}}{3H}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)+\frac{2i\omega_{0}^{\pm}}{9T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)\\ \times\left[\frac{6\sigma^{\omega}_{A}}{T^{2}}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)\omega_{0}^{\pm}+\frac{i}{\tau_{A}\tau_{H}}\right]=0.$ (116)
Expanding $\omega_{0}$ in powers of $T/\lvert\mu_{V}\rvert$ gives, to leading order in $T/\lvert\mu_{V}\rvert$:
$\omega_{\pm}\simeq\dfrac{-i\pi^{2}T^{2}}{6(\tau_{A}+\tau_{H})\mu_{V}^{2}}\\ \times\left(1\pm\sqrt{1-\frac{48}{\pi^{2}}(\tau_{A}+\tau_{H})^{2}\kappa_{\Omega}^{2}}\right).$ (117)
Note that when both relaxation times are finite, there is no propagation for small wavenumbers. This is in contrast to the result obtained in Subsec. 3.3, where we considered $\tau_{A}=\infty$, and the analogous modes propagate for all wavelengths. Notably, the $\pm$ modes are sensitive to an emergent effective relaxation time $\tau_{\rm eff}=\tau_{A}+\tau_{H}$ which can be traced to the interdependence of the axial and helical degrees of freedom in the degenerate limit. From Eq. (117), we extract the critical wavenumber $k_{*}$ above which these two modes start propagating:
$k_{*}=\dfrac{\kappa_{\Omega}^{*}T}{\Omega}=\frac{\pi}{4\sqrt{3}}\frac{T}{(\tau_{A}+\tau_{H})\Omega}\,.$ (118)
@@
$\delta\mu^{*}_{\chi}=\kappa_{\Omega}\delta\mu^{*}_{V}\times$
$\dfrac{\frac{2\omega_{*}}{T^{2}}(A+s_{V}B)(\sigma^{\omega}_{A}-s_{V}\sigma^{\omega}_{H})+\frac{i}{3}\left(\frac{A}{\tau_{H}}+\frac{s_{V}B}{\tau_{A}}\right)}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}},$
$\delta\mu^{*}_{\tilde{\chi}}=\kappa_{\Omega}\delta\mu^{*}_{V}\times$
$\dfrac{\frac{2\omega_{*}}{T^{2}}(A-s_{V}B)(\sigma^{\omega}_{A}+s_{V}\sigma^{\omega}_{H})+\frac{i}{3}\left(\frac{A}{\tau_{H}}-\frac{s_{V}B}{\tau_{A}}\right)}{\left(\frac{2\omega_{*}}{T^{2}}\right)^{2}[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]+\frac{2i\omega_{*}}{3T^{2}}\sigma^{\omega}_{A}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)-\frac{1}{9\tau_{A}\tau_{H}}}.$ (119)
@@
$\delta\mu_{\chi}^{\pm}$ $\simeq\frac{AT^{2}}{4\sigma_{A}^{\omega}}\frac{\kappa_{\Omega}(\tau_{A}+\tau_{H})}{\omega_{\pm}(\tau_{A}+\tau_{H})+\frac{iT^{2}}{6\sigma^{\omega}_{A}}}\delta\mu_{V}^{\pm}$
$\simeq-\frac{\kappa_{\Omega}(\tau_{A}+\tau_{H})}{3\alpha_{V}^{2}\omega_{\mp}(\tau_{A}+\tau_{H})}\dfrac{\lvert\mu_{V}\rvert}{T}\delta\mu_{V}^{\pm},$
$\delta\mu_{\tilde{\chi}}^{\pm}$ $\simeq\frac{AT^{2}}{4\sigma_{A}^{\omega}}\frac{\kappa_{\Omega}(\tau_{A}-\tau_{H})}{\omega_{\pm}(\tau_{A}+\tau_{H})+\frac{iT^{2}}{6\sigma_{A}}}\delta\mu_{V}^{\pm}$
$\simeq-\dfrac{\kappa_{\Omega}(\tau_{A}-\tau_{H})}{3\alpha_{V}^{2}\omega_{\mp}(\tau_{A}+\tau_{H})}\dfrac{\lvert\mu_{V}\rvert}{T}\delta\mu_{V}^{\pm}.$ (120)
As in the case with $\tau_{A}\rightarrow\infty$ considered in Eq. (78), the amplitudes of $\delta\mu^{\pm}_{\chi}$ and $\delta\mu^{\pm}_{\tilde{\chi}}$ are $\lvert\alpha_{V}\rvert=\lvert\mu_{V}\rvert/T$ times larger than $\delta\mu_{V}^{\pm}$, due to the behaviour $\omega_{\mp}\sim\alpha_{V}^{-2}$. Contrary to Eq. (78), the degeneracy between $\delta\mu^{\pm}_{\chi}$ and $\delta\mu^{\pm}_{\tilde{\chi}}$ is lifted whenever both $\tau_{A}$ and $\tau_{H}$ are finite (i.e., non-zero and non-infinite).
@@
$\delta\mu_{\chi}^{\textrm{im.}}\simeq-\dfrac{2(\tau_{A}-\tau_{H})}{(\tau_{A}+\tau_{H})\alpha_{V}^{2}}e^{-\lvert\alpha_{V}\rvert}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}\,,$
$\delta\mu_{V}^{\textrm{im.}}\simeq-\dfrac{72i\kappa_{\Omega}\tau_{A}\tau_{H}}{\pi^{2}(\tau_{A}+\tau_{H})\alpha_{V}^{2}}e^{-2\lvert\alpha_{V}\rvert}\delta\mu_{\tilde{\chi}}^{\textrm{im.}}$ (121)
We present in Fig. 7 the propagation properties of an initial Gaussian fluctuation in the vector chemical potential, given in Eq. (57), similar to the one studied in Fig. 3 ($\alpha_{V}=20$, $\tau_{H}=2\sigma/\tilde{v}$; see the end of Sec. 3.3 for details). As opposed to Fig. 3, here we take a finite axial relaxation time, which we set to $\tau_{A}=\tau_{H}/2$. While the propagation of $\delta\bar{\mu}_{V}(t,z)$ and of $\delta\bar{\mu}_{\chi}(t,z)$ remains unchanged, the amplitude of $\delta\bar{\mu}_{\tilde{\chi}}(t,z)$ is reduced by a factor of about $3$ compared to the case when the axial charge is conserved.
@@ 4.6 Generic properties of spectra
Figure 8: Solution of Eq. (127) showing the values of the parameters (blue curve) for which the spectrum of excitations contains a merging point. The curve divides the parameter space into two regions: Above the curve, the spectrum has only one branch of propagating modes; Below the curve, the spectrum displays two disconnected branches of propagating modes.
@@
In Sec. 3.4, we have seen that there exists a value of the vector chemical potential which divides the spectrum of excitations into two regions, depending on whether there exist one or two branches of propagating modes (c.f. Fig. 5). The chemical potential separating both regions was found to be $\mu_{V}=1.4471\,T$, and the wavenumber at which the two propagating branches merge satisfies $\tau_{H}\kappa_{\Omega}=3.0642$. In this section, the presence of a finite axial relaxation time $\tau_{A}$ provides an extra parameter, and we expect to find a one-parameter family of such points. We will impose the existence of the merging point in order to find the curve separating both regions in terms of $\mu_{V}$ and $\tau_{H}/\tau_{A}\,$.
@@
$a(\alpha_{V})\omega^{3}+i\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)b(\alpha_{V})\omega^{2}-$
$\left(c(\alpha_{V})\kappa_{\Omega}^{2}+\dfrac{\varsigma(\alpha_{V})}{\tau_{A}\tau_{H}}\right)\omega-i\kappa_{\Omega}^{2}\delta(\alpha_{V},\tau_{A},\tau_{H})=0\,,$ (122)
where $a\,,b$ and $c$ are already given in (73), while $\varsigma$ and $\delta$ are defined as
$\varsigma=\dfrac{2}{9T^{2}}\left(\sigma_{A}^{\omega}-\frac{1}{3}T^{2}\Delta H\right)\,,\ \ \ \delta=\dfrac{A^{2}\tau_{A}+B^{2}\tau_{H}}{3H\tau_{A}\tau_{H}}\,.$ (123)
@@
$3a\omega_{\textrm{merg.}}^{2}+2ib\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)\omega_{\textrm{merg.}}$
$-\left(c\kappa_{\textrm{merg.}}^{2}+\dfrac{\varsigma}{\tau_{A}\tau_{H}}\right)$ $=0\,,$
@@
$\omega_{\textrm{merg.}}$ $=-\dfrac{ib}{3a}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)\,$
$\kappa_{\textrm{merg.}}^{2}$ $=\dfrac{b^{2}}{3ac}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)^{2}-\dfrac{\varsigma}{c}\frac{1}{\tau_{A}\tau_{H}}\,.$ (125)